# The Tarantula Massive Binary Monitoring

### v1

@@ 5 Discussion and conclusion
Based on our disentangling analysis, we discuss here the potential nature of the system and prospects to establish or reject VFTS 812 as an O+BH binary. We obtain a detection threshold of for MS companions, with lower mass companions avoiding detection. The orbital constraint on the other hand implies at 5% significance. Subsequently, we consider a few possible scenarios:
@@
MS companion: It is possible that the companion is an MS star with a mass between 5.1 and 6 , which satisfies the orbital constraint yet remains elusive. We are limited by current observations in testing the presence of such a companion. However, in the scenario where VFTS 812 is formed by two young MS stars, its location in 30 Dor remains intriguing: VFTS 812 is offset by 4.9′ from R136, the star forming region compatible with the age of VFTS 812 in the 30 Dor complex (Schneider et al., 2018). VFTS 812 is also noticeably isolated; the few stars in its neighborhood are much older B-type stars (Schneider et al., 2018). Furthermore, it shows no indication of a runaway motion, whether in its systemic radial velocity (Sana et al., 2022), or in its Gaia proper motion (Stoop et al., 2024), invalidating a scenario where it has been dynamically ejected from R136.
@@
Hot stripped companion: A hot stripped companion formed via binary mass transfer with peaks in the ultraviolet (UV) regime and is considerably smaller in radius compared to similar-mass MS stars. It thus contributes even less in optical wavelengths in comparison. An excess flux in the ultraviolet (UV) is an effective way to detect stripped companions (Ludwig et al., 2025); however, the UV excess contribution of a stripped star is expected to vanish next to a MS companion (see Section S2.1 in Drout et al., 2023). The photometric AB-magnitude colors of VFTS 812 using Swift and Bessel filters: UVM2 – V = 0.78, UVW1 – B = 0.64, U – V = 0.01 are all far from passing the UV excess criteria of Ludwig et al. (2025). For companions with where stripped stars transition to the Wolf-Rayet regime (Shenar et al., 2019), the companion would be much more luminous and detectable. On the other hand, for , factors such as an extreme mass ratio, high eccentricity, no evidence of high rotational velocity and He or N enrichment in the primary pose significant challenges for this scenario. We therefore consider the presence of a hot stripped companion unlikely.
@@
Compact companion: If the companion is a compact object, the orbital constraint on essentially requires a BH companion. A significant kick during BH formation could explain its isolated location and eccentricity (Mahy et al., 2022; Willcox et al., 2025). Although the primary lacks obvious signatures of past interaction, if we consider that it was rejuvenated via accretion followed by BH formation with a kick, we can explain its young age, isolated location and eccentricity.
@@
Fast rotating primary: In Section 3.2, we excluded high rotational velocity () solutions for the primary star from BONNSAI. If the primary was a fast rotator, the following can be said about potential companions: (i) For an MS or stripped companion with the primary rotational axis and the orbital axis aligned, the orbital inclination would be . This would imply a significantly higher minimum companion mass, favoring its detection. (ii) For a BH companion, a BH formation kick could disturb the above alignment and similar constraints on its mass cannot be made, although the remaining arguments made earlier about a BH companion would still apply.
@@
Given these formation scenario challenges, VFTS 812 remains a compelling testbed for binary evolution models. The major limitation of our data is the S/N, which holds the key to confirm the presence or absence of a luminous companion. We encourage follow-up high-S/N spectroscopic observations to confirm the nature of the unseen companion in VFTS 812. Complementary far-ultraviolet spectroscopy can also greatly assist in determining accurate atmospheric parameters of the primary.
@@ Acknowledgements.
This research received support from the Flemish Government under the long-term structural Methusalem funding program, project SOUL: Stellar evolution in full glory, grant METH/24/012 at KU Leuven. SJ acknowledges support from the Japan Society for the Promotion of Science (JSPS) as a postdoctoral fellow. JIV acknowledges support from the European Research Council for the ERC Advanced Grant 101054731. TS acknowledges support from the Israel Science Foundation (ISF) under grant number 0603225041 and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement 101164755/METAL).References
- Clues on the Origin and Evolution of Massive Contact Binaries: Atmosphere Analysis of VFTS 352. ApJ 880 (2), pp. 115. External Links: Document, 1906.01066, ADS entry Cited by: Appendix C.
- The Tarantula Massive Binary Monitoring. I. Observational campaign and OB-type spectroscopic binaries. A&A 598, pp. A84. External Links: Document, 1610.03500, ADS entry Cited by: §D.1, §2, §4.1.
- The Chemical Composition of the Sun. ARA&A 47 (1), pp. 481–522. External Links: Document, 0909.0948, ADS entry Cited by: Appendix C.
- X-Shooting ULLYSES: Massive stars at low metallicity: XII. Clumped winds of O-type (super)giants in the Large Magellanic Cloud. A&A 697, pp. A54. External Links: Document, 2503.14687, ADS entry Cited by: Appendix C, §3.1.
- The R136 star cluster dissected with Hubble Space Telescope/STIS. III. The most massive stars and their clumped winds. A&A 663, pp. A36. External Links: Document, 2202.11080, ADS entry Cited by: Appendix C, §3.1.
- Rotating massive main-sequence stars. I. Grids of evolutionary models and isochrones. A&A 530, pp. A115. External Links: Document, 1102.0530, ADS entry Cited by: §3.2.
- 2D unified atmosphere and wind simulations of O-type stars. A&A 684, pp. A177. External Links: Document, 2401.08391, ADS entry Cited by: Appendix C.
- An observed population of intermediate-mass helium stars that have been stripped in binaries. Science 382 (6676), pp. 1287–1291. External Links: Document, 2307.00061, ADS entry Cited by: §5.
- The VLT-FLAMES Tarantula Survey. I. Introduction and observational overview. A&A 530, pp. A108. External Links: Document, 1103.5386, ADS entry Cited by: §2.
- spinOS: SPectroscopic and INterferometric Orbital Solution finder. Note: Astrophysics Source Code Library, record ascl:2102.001 External Links: ADS entry Cited by: §D.1.
- Correcting for the Effects of Interstellar Extinction. PASP 111 (755), pp. 63–75. External Links: Document, astro-ph/9809387, ADS entry Cited by: Appendix C.
- Improving 1D stellar atmosphere models with insights from multi-dimensional simulations: I. 1D versus 2D stratifications and spectral comparisons for O stars. A&A 694, pp. A269. External Links: Document, 2501.14511, ADS entry Cited by: Appendix C, §3.1.
- Orbital elements of multiple spectroscopic stars.. A&AS 114, pp. 393. External Links: ADS entry Cited by: §1.
- Empirical mass-loss rates and clumping properties of Galactic early-type O supergiants. A&A 655, pp. A67. External Links: Document, 2108.08340, ADS entry Cited by: Appendix C.
- The mass and helium discrepancy in massive young stars. In The Atmospheres of Early-Type Stars, U. Heber and C. S. Jeffery (Eds.), Vol. 401, pp. 21. External Links: Document, ADS entry Cited by: §3.1.
- Uncovering astrometric black hole binaries with massive main-sequence companions with Gaia. A&A 658, pp. A129. External Links: Document, 2111.06427, ADS entry Cited by: §1.
- Properties of OB star-black hole systems derived from detailed binary evolution models. A&A 638, pp. A39. External Links: Document, 1912.09826, ADS entry Cited by: §1.
- The Stripped-Star Ultraviolet Magellanic Cloud Survey (SUMS): The UV Photometric Catalog and Stripped Star Candidate Selection. ApJ, pp. in press. External Links: Document, 2505.18632, ADS entry Cited by: §5.
- Identifying quiescent compact objects in massive Galactic single-lined spectroscopic binaries. A&A 664, pp. A159. External Links: Document, 2207.07752, ADS entry Cited by: §1, §5.
- 2D unified atmosphere and wind simulations for a grid of O-type stars. A&A 704, pp. A121. External Links: Document, 2511.04542, ADS entry Cited by: Appendix C, §3.1.
- Spectral analysis of early-type stars using a genetic algorithm based fitting method. A&A 441 (2), pp. 711–733. External Links: Document, astro-ph/0506751, ADS entry Cited by: Appendix C.
- Realistic Predictions for Gaia Black Hole Discoveries: Comparison of Isolated Binary and Dynamical Formation Models. PASP 137 (4), pp. 044202. External Links: Document, 2502.03527, ADS entry Cited by: §1.
- A distance to the Large Magellanic Cloud that is precise to one per cent. Nature 567 (7747), pp. 200–203. External Links: Document, 1903.08096, ADS entry Cited by: Appendix C.
- Atmospheric NLTE-models for the spectroscopic analysis of blue stars with winds. II. Line-blanketed models. A&A 435 (2), pp. 669–698. External Links: Document, astro-ph/0411398, ADS entry Cited by: Appendix C.
- The VLT-FLAMES Tarantula Survey. VIII. Multiplicity properties of the O-type star population. A&A 550, pp. A107. External Links: Document, 1209.4638, ADS entry Cited by: §D.1.
- The VLT-FLAMES Tarantula Survey. Observational evidence for two distinct populations of massive runaway stars in 30 Doradus. A&A 668, pp. L5. External Links: Document, 2211.13476, ADS entry Cited by: §5.
- Atmospheric NLTE-models for the spectroscopic analysis of luminous blue stars with winds.. A&A 323, pp. 488–512. External Links: ADS entry Cited by: Appendix C, §3.1.
- Bonnsai: a Bayesian tool for comparing stars with stellar evolution models. A&A 570, pp. A66. External Links: Document, 1408.3409, ADS entry Cited by: §3.2.
- The VLT-FLAMES Tarantula Survey. XXIX. Massive star formation in the local 30 Doradus starburst. A&A 618, pp. A73. External Links: Document, 1807.03821, ADS entry Cited by: §3, §5.
- The Wolf-Rayet binaries of the nitrogen sequence in the Large Magellanic Cloud. Spectroscopy, orbital analysis, formation, and evolution. A&A 627, pp. A151. External Links: Document, 1905.09296, ADS entry Cited by: §5.
- The Tarantula Massive Binary Monitoring. VI. Characterisation of hidden companions in 51 single-lined O-type binaries: A flat mass-ratio distribution and black-hole binary candidates. A&A 665, pp. A148. External Links: Document, 2207.07674, ADS entry Cited by: Appendix D, Appendix D, Appendix E, §1, §1, §3.1, §3.3, §3.3, §3, §4.1, §4.
- An X-ray-quiet black hole born with a negligible kick in a massive binary within the Large Magellanic Cloud. Nature Astronomy 6, pp. 1085–1092. External Links: Document, 2207.07675, ADS entry Cited by: §D.2, §1, §1.
- Two waves of massive stars running away from the young cluster R136. Nature 634 (8035), pp. 809–812. External Links: Document, 2410.06255, ADS entry Cited by: §5.
- Atmospheric NLTE models for the spectroscopic analysis of blue stars with winds. IV. Porosity in physical and velocity space. A&A 619, pp. A59. External Links: Document, 1805.11010, ADS entry Cited by: Appendix C, §3.1.
- The properties of ten O-type stars in the low-metallicity galaxies IC 1613, WLM, and NGC 3109. A&A 572, pp. A36. External Links: Document, 1410.0934, ADS entry Cited by: Appendix C.
- Cosmic-Ray Rejection by Laplacian Edge Detection. PASP 113 (789), pp. 1420–1427. External Links: Document, astro-ph/0108003, ADS entry Cited by: §2.
- X-Shooting ULLYSES: Massive Stars at low metallicity: IX. Empirical constraints on mass-loss rates and clumping parameters for OB supergiants in the Large Magellanic Cloud. A&A 692, pp. A91. External Links: Document, 2410.14937, ADS entry Cited by: §3.1.
- Constraints on Neutrino Natal Kicks from Black-Hole Binary VFTS 243. Phys. Rev. Lett. 132 (19), pp. 191403. External Links: Document, 2310.01509, ADS entry Cited by: §1.
- Binarity at LOw Metallicity (BLOeM): Bayesian inference of natal kicks from inert black hole binaries. A&A 700, pp. A59. External Links: Document, 2504.16669, ADS entry Cited by: §1, §5.
Appendix A Summary of observations
@@
Notes: For every epoch, we provide the MJD at the start of the observation (column 2), grade assigned by ESO (column 3), and S/N measured in the continuum region between 4050 – 4060 (column 4). For epochs marked with an asterisk, the data quality was too poor to extract spectra.
Appendix B Optimal sky subtraction
@@
The spatial and temporal variation in the Tarantula nebula makes it difficult to obtain perfect sky subtraction. The IFU approach is helpful to remedy this by capturing the sky background in the vicinity of the star for every epoch. However, we noticed two important features in our observations: (i) The star is not always aligned on the central four spaxels and changes positions on the IFU from epoch to epoch (Fig. B.1); and (ii) The nebular emission in the immediate vicinity of the star is not always identical to that in the star itself (Fig. B.2). Perfect sky subtraction is thus not guaranteed even with IFU observations. This was considered when choosing the spaxels for the star and sky regions. Here, we discuss two approaches along with their suitability for spectral disentangling.
@@
In the first approach, we prioritized perfect nebular subtraction to leverage the capabilities of the IFU mode. For each epoch, we selected the four brightest spaxels for the star and summed them up. To obtain perfect subtraction, we tried different combinations of four other spaxels to sum up and find a suitable sky spectrum. This was done for each epoch individually, leading to varying choices for sky spaxels across epochs. In many cases, we chose sky spaxels immediately next to the star spaxels for the best possible subtraction. This came at the cost of subtracting some of the star signal present in these spaxels, reducing the signal-to-noise ratio (S/N) of the resulting spectrum. Despite these considerations, the nebular subtraction in many spectra remained imperfect and variable across epochs as seen in Figure B.2. This can severely hamper spectral disentangling when dealing with faint/unseen companions that require the best quality data to obtain meaningful constraints. Furthermore, the compromise on S/N is yet another factor making disentangling more challenging.
@@
In the second approach, given that the nebular subtraction leaving small residuals was inevitable, we prioritized S/N which is particularly important when searching for a faint companion. We followed the same procedure as before to select the four star spaxels. However, for the sky subtraction, we chose four spaxels farthest from the star that were unlikely to have any significant signal from the star. This resulted in a more accurate sky subtraction for the continuum region, while the nebular subtraction was slightly compromised. To further optimize the process, we took advantage of the fact that the sky background did not have any spectral lines in our wavelength range other than the Hδ nebular emission line. We divided the sky spectrum into two parts: the Hδ line spanning the small wavelength range of 4103.9 – 4106.7 , and the continuum region covering the remaining wavelength range. The latter was modeled with a cubic polynomial and combined with actual data within the Hδ line to form a new sky spectrum. The main motivation behind doing this was to avoid adding extra continuum noise in the final sky subtracted spectra. Although the nebular subtraction in this case slightly worsened compared to the previous approach, it is circumvented successfully in our modified disentangling method discussed in Section 4.
Appendix C Atmospheric analysis
@@
We use fastwind (Santolaya-Rey et al. 1997; Sundqvist and Puls 2018) and Kiwi-GA (Brands et al. 2022, 2025) to fit the spectral lines of VFTS 812. This method is an evolution from previous genetic algorithm based fitting solutions (Mokiem et al. 2005; Tramper et al. 2014; Abdul-Masih et al. 2019; Hawcroft et al. 2021). We allow for 6 fit parameters (, , , , , and nitrogen abundance). Here is the only broadening mechanism included meaning this value could be read as the combination of macro turbulence and rotational broadening. The micro turbulence is kept fixed at 10 km s-1. The nitrogen number abundance N is expressed in . We used fastwind version 10.5 (Puls et al. 2005) with metallicity scaled to 0.5 times the solar metallicity (Asplund et al. 2009) and a smooth wind. In total we have access to 17 spectral windows (Figure C.1) containing a total of 41 diagnostic spectral lines (Table 2). In this method all lines were fitted at the same time and the total goodness of fit was taken into account to decide the quality of the fit. This means that even lines which are generally not thought to be indicative lines of a parameter still contribute to the determination of that parameter. To determine the radius of the star, we used the observed K-band magnitude of 14.17, an estimated 50 kpc distance to the LMC (Pietrzyński et al. 2019), and the extinction law Fitzpatrick (1999) to scale the radius such that the model spectral energy distribution crosses the K-band luminosity. Table 3 lists the best fit values with their 1- uncertainty.
@@
When using the unaltered version of fastwind, turbulent pressure is not included into the calculation of the stellar structure. Debnath et al. (2024) showed that the 2D average stellar structure is very different from the 1D models used for spectral synthesis. Additionally, they showed that by including turbulent pressure in the hydrostatic equation, it is possible to offset this difference. Moens et al. (2025) showed through a grid of multi-D models that the turbulent pressure is highly dependent on the Eddington factor of the star, giving us an indication of which turbulent pressure to include. Subsequently, the derived surface gravity, and the spectroscopic mass goes up as the turbulent pressure increases (González-Torà et al. 2025, Verhamme et al. in prep). All this led us to try a spectral fit in which we included a constant turbulent velocity of 50 km s-1. The results of this fit did not change any parameters outside of the error-margins besides from the surface gravity and thus the resulting derived spectroscopic mass. The results of this fit are also given in Table 3 next to the normal fit for ease of comparison. Figure C.1 shows the best fit models with and without turbulent pressure, compared to the data with the 1- range of the spectra. The spectroscopic mass strongly depends on the chosen turbulent velocity, which is only loosely bound by the radiation hydrodynamical models (Moens et al. 2025). Because it is not possible to constrain the turbulent velocity observationally, strong constraints on spectroscopic mass are inaccessible for VFTS 812.
@@
Notes: The first column shows the atom and its ionization stage which is responsible for the transition. The second column shows the corresponding rest wavelength with possible multiplets. The third column shows where to find this line in the fit summary (Fig. C.1).
@@
Appendix D Spectral disentangling
A comprehensive description of the shift-and-technique is provided in Section 3.1 of TS22. Making an agnostic assumption about the binary being an SB2, a 2-dimensional grid disentangling can be used to constrain the respective RV semi-amplitudes and of the primary and secondary components. A chi-squared minimization can yield the best combination, where the chi-squared is calculated as

### v2

@@ 5 Discussion and conclusion
Based on our disentangling analysis, we discuss here the potential nature of the system and prospects for establishing or rejecting VFTS 812 as an O+BH binary. We obtained a detection threshold of for MS companions (i.e., lower-mass companions avoid detection). The orbital constraint, on the other hand, implies at 5% significance. Subsequently, we considered a few possible scenarios:
@@
MS companion: It is possible that the companion is an MS star with a mass between 5.1 and 6 . This satisfies the orbital constraint, but any such star remains elusive. Our means for detecting such a companion are limited by current observations. However, in the scenario where VFTS 812 is formed by two young MS stars, its location in 30 Dor remains intriguing: VFTS 812 is offset by 4.9′ from R136, the star-forming region compatible with the age of VFTS 812 in the 30 Dor complex (Schneider et al., 2018). VFTS 812 is also noticeably isolated; the few stars in its neighborhood are much older B-type stars (Schneider et al., 2018). Furthermore, it shows no indication of a runaway motion, neither in its systemic RV (Sana et al., 2022) nor in its Gaia proper motion (Stoop et al., 2024), meaning it could not have been dynamically ejected from R136.
@@
Hot stripped companion: A hot stripped companion formed via binary mass transfer with peaks in the ultraviolet (UV) regime and is considerably smaller in radius compared to similar-mass MS stars. It thus contributes even less in optical wavelengths in comparison. An excess flux in the UV is an effective way to detect stripped companions (Ludwig et al., 2025); however, the UV excess contribution of a stripped star is expected to vanish next to a MS companion (see Sect. S2.1 in Drout et al., 2023). The photometric AB-magnitude colors of VFTS 812 using Swift and Bessel filters (UVM2 – V = 0.78, UVW1 – B = 0.64, U – V = 0.01) are all far from passing the UV excess criteria of Ludwig et al. (2025). For companions with where stripped stars transition to the Wolf-Rayet regime (Shenar et al., 2019), the companion would be much more luminous and detectable. On the other hand, for , factors such as an extreme mass ratio, high eccentricity, no evidence of high rotational velocity, and He or N enrichment in the primary pose significant challenges for this scenario. We therefore consider the presence of a hot stripped companion unlikely.
@@
Compact companion: If the companion is a compact object, the orbital constraint on essentially requires a BH companion. A significant kick during BH formation could explain its isolated location and eccentricity (Mahy et al., 2022; Willcox et al., 2025). Although the primary lacks obvious signatures of a past interaction, if we consider that it was rejuvenated via accretion followed by BH formation with a kick, we can explain its young age, isolated location, and eccentricity.
@@
Fast rotating primary: In Sect. 3.2 we excluded high rotational velocity () solutions for the primary star from BONNSAI. If the primary is a fast rotator, the following can be said about potential companions: (i) For an MS or stripped companion whose primary rotational axis and orbital axis are aligned, the orbital inclination would be . This would imply a significantly higher minimum companion mass and a high probability of being detectable. (ii) For a BH companion, a BH formation kick could disturb the above alignment, and similar constraints on its mass could not be made, although the remaining arguments made earlier about a BH companion would still apply.
@@
Given these formation scenario challenges, VFTS 812 remains a compelling testbed for binary evolution models. The major limitation of our data is the S/N, which holds the key to confirm the presence or absence of a luminous companion. We encourage follow-up high-S/N spectroscopic observations to confirm the nature of the unseen companion in VFTS 812. Complementary far-UV spectroscopy can also greatly assist in determining accurate atmospheric parameters for the primary.
Appendix A Acknowledgements
@@ Acknowledgements.
This research received support from the Flemish Government under the long-term structural Methusalem funding program, project SOUL: Stellar evolution in full glory, grant METH/24/012 at KU Leuven. SJ acknowledges support from the Japan Society for the Promotion of Science (JSPS) as a postdoctoral fellow. JIV acknowledges support from the European Research Council for the ERC Advanced Grant 101054731. TS acknowledges support from the Israel Science Foundation (ISF) under grant number 0603225041 and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement 101164755/METAL).Appendix B Summary of observations
@@
Appendix C Optimal sky subtraction
@@
The spatial and temporal variation in the Tarantula nebula makes it difficult to obtain perfect sky subtraction. The IFU approach is helpful to remedy this by capturing the sky background in the vicinity of the star for every epoch. However, we noticed two important features in our observations: (i) The star is not always aligned on the central four spaxels and changes positions on the IFU from epoch to epoch (Fig. C.1); and (ii) The nebular emission in the immediate vicinity of the star is not always identical to that in the star itself (Fig. C.2). Perfect sky subtraction is thus not guaranteed even with IFU observations. This was considered when choosing the spaxels for the star and sky regions. Here, we discuss two approaches along with their suitability for spectral disentangling.
@@
In the first approach, we prioritized perfect nebular subtraction to leverage the capabilities of the IFU mode. For each epoch, we selected the four brightest spaxels for the star and summed them up. To obtain perfect subtraction, we tried different combinations of four other spaxels to sum up and find a suitable sky spectrum. This was done for each epoch individually, leading to varying choices for sky spaxels across epochs. In many cases, we chose sky spaxels immediately next to the star spaxels for the best possible subtraction. This came at the cost of subtracting some of the star signal present in these spaxels, reducing the signal-to-noise ratio (S/N) of the resulting spectrum. Despite these considerations, the nebular subtraction in many spectra remained imperfect and variable across epochs as seen in Fig. C.2. This can severely hamper spectral disentangling when dealing with faint/unseen companions that require the best quality data to obtain meaningful constraints. Furthermore, the compromise on S/N is yet another factor making disentangling more challenging.
@@
In the second approach, given that the nebular subtraction leaving small residuals was inevitable, we prioritized S/N which is particularly important when searching for a faint companion. We followed the same procedure as before to select the four star spaxels. However, for the sky subtraction, we chose four spaxels farthest from the star that were unlikely to have any significant signal from the star. This resulted in a more accurate sky subtraction for the continuum region, while the nebular subtraction was slightly compromised. To further optimize the process, we took advantage of the fact that the sky background did not have any spectral lines in our wavelength range other than the Hδ nebular emission line. We divided the sky spectrum into two parts: the Hδ line spanning the small wavelength range of 4103.9 – 4106.7 , and the continuum region covering the remaining wavelength range. The latter was modeled with a cubic polynomial and combined with actual data within the Hδ line to form a new sky spectrum. The main motivation behind doing this was to avoid adding extra continuum noise in the final sky subtracted spectra. Although the nebular subtraction in this case slightly worsened compared to the previous approach, it is circumvented successfully in our modified disentangling method discussed in Sect. 4.
Appendix D Atmospheric analysis
@@
We use fastwind (Santolaya-Rey et al. 1997; Sundqvist and Puls 2018) and Kiwi-GA (Brands et al. 2022, 2025) to fit the spectral lines of VFTS 812. This method is an evolution from previous genetic algorithm based fitting solutions (Mokiem et al. 2005; Tramper et al. 2014; Abdul-Masih et al. 2019; Hawcroft et al. 2021). We allow for 6 fit parameters (, , , , , and nitrogen abundance). Here is the only broadening mechanism included meaning this value could be read as the combination of macro turbulence and rotational broadening. The micro turbulence is kept fixed at 10 km s-1. The nitrogen number abundance N is expressed in . We used fastwind version 10.5 (Puls et al. 2005) with metallicity scaled to 0.5 times the solar metallicity (Asplund et al. 2009) and a smooth wind. In total we have access to 17 spectral windows (Fig. D.1) containing a total of 41 diagnostic spectral lines (Table 2). In this method all lines were fitted at the same time and the total goodness of fit was taken into account to decide the quality of the fit. This means that even lines which are generally not thought to be indicative lines of a parameter still contribute to the determination of that parameter. To determine the radius of the star, we used the observed K-band magnitude of 14.17, an estimated 50 kpc distance to the Large Magellanic Cloud (Pietrzyński et al. 2019), and the extinction law Fitzpatrick (1999) to scale the radius such that the model spectral energy distribution crosses the K-band luminosity. Table 3 lists the best-fit values with their 1 uncertainty.
@@
When using the unaltered version of fastwind, turbulent pressure is not included into the calculation of the stellar structure. Debnath et al. (2024) showed that the 2D average stellar structure is very different from the 1D models used for spectral synthesis. Additionally, they showed that by including turbulent pressure in the hydrostatic equation, it is possible to offset this difference. Moens et al. (2025) showed through a grid of multi-D models that the turbulent pressure is highly dependent on the Eddington factor of the star, giving us an indication of which turbulent pressure to include. Subsequently, the derived surface gravity, and the spectroscopic mass goes up as the turbulent pressure increases (González-Torà et al. 2025, Verhamme et al. in prep). All this led us to try a spectral fit in which we included a constant turbulent velocity of 50 km s-1. The results of this fit did not change any parameters outside of the error-margins besides from the surface gravity and thus the resulting derived spectroscopic mass. The results of this fit are also given in Table 3 next to the normal fit for ease of comparison. Figure D.1 shows the best-fit models with and without turbulent pressure, compared to the data with the 1 range of the spectra. The spectroscopic mass strongly depends on the chosen turbulent velocity, which is only loosely bound by the radiation hydrodynamical models (Moens et al. 2025). Because it is not possible to constrain the turbulent velocity observationally, strong constraints on spectroscopic mass are inaccessible for VFTS 812.
@@
Appendix E Photometric constraints
In Sect. 3.3 the lack of observed photometric variability in VFTS 812 was used to obtain an upper limit on the inclination, on the basis that a more edge-on orbit would produce an observable primary eclipse. Figure E.1 shows the OGLE I-band light curve that TS22 remarked on for not having any significant frequencies. Using a 3 threshold, we find that the flux-dip from an eclipse has to be 2.6% for detection. From spectral disentangling, we know that the flux contribution of the secondary is 1.6% (see Sect. 4.2). A full eclipse of the primary by a companion with a radius of will cover 5% of the primary surface, while contributing 1.6% to the total flux, thus resulting in a flux-dip of 3.4%. While partial eclipses can be considered for finer estimation of eclipse detectability, we adopt the full eclipse scenario to remain conservative. The lack of a detected eclipse therefore enables an additional inclination constraint on the MS companion scenario.
Appendix F Spectral disentangling
A comprehensive description of the shift-and-technique is provided in Sect. 3.1 of TS22. Making an agnostic assumption about the binary being an SB2, a 2D grid disentangling can be used to constrain the respective RV semi-amplitudes and of the primary and secondary components. A chi-squared minimization can yield the best combination, where the chi-squared is calculated as